OpenStudy (anonymous):

There are 650 students. 44% Plan to attend graduation and 48% plan to skip. The rest hasnt decided. What % hasnt decided?

OpenStudy (oaktree):

8%.

OpenStudy (oaktree):

This is because 44+48=92.

OpenStudy (oaktree):

100-92=8, so the answer is 8%.

OpenStudy (anonymous):

Ok next Q . 8% = How many students?

OpenStudy (oaktree):

Multiply (0.08) by (650) and you will have your answer.

OpenStudy (oaktree):

In this case, the answer is 52.

OpenStudy (anonymous):

52?

OpenStudy (oaktree):

Yup.

OpenStudy (diyadiya):

650*44/100=286 studnts 650*48/100=312 students total = 598 650-598=52

OpenStudy (oaktree):

Diyadiya's way works too, but it takes more time.

OpenStudy (anonymous):

oh haha you answered right before I got to lol so I was asking if it was 52 and you got it right befor me haha (: Thanks for the help!!!!!

OpenStudy (oaktree):

No problem!